# Causality for Earth Science - A Review on Time-series and Spatiotemporal Causality Methods

### v1

@@ 5.2 Realistic Simulated Data
The realistic simulated Functional Magnetic Resonance Imaging dataset contains BOLD (Blood-oxygen-level dependent) datasets for 28 different simulations for different brain networks (Smith et al., 2011). Each dataset contains the neural activity, based on the blood flow change. Each dataset contains up to 50 time series which contains between 50 and 5000 time points. Note that these datasets are considered causally sufficient.
DREAM4:
(Marbach et al., 2009) is a synthetic gene expression dataset, generated by software, and based on patterns found in model organisms. This dataset consisted of both time-series and steady-state data and the objective is to learn gene regulatory networks from gene expression data. DREAM4 consists of 5 independent datasets, each with 10 different time-series recordings for 100 genes across 21 time steps.
NCEP-DOE Reanalysis 2 product:
NCEP-DOE Reanalysis 2 product (Kanamitsu et al., 2002) for the time period 1979 to current, is a data product provided by the US National Centers for Environmental Protection (NCEP) and the Department of Energy (DOE). Reanalysis data sets are generated by fitting a complex climate model to all available data for a given period of time, thus generating estimates for times and locations that were not originally observed. A large set of almost 40 atmospheric variables are measured in the reanalysis dataset. The dataset covers 90N - 90S, 0E - 357.5E with a 2.5-degree latitude x 2.5-degree longitude global grid (144x73).
CausalWorld:
(Ahmed et al., 2021) is an open-source benchmarking for causal structure learning. This platform contains a robotic environment manipulation dataset for different tasks. These datasets represent different causal structures of interacting objects like robot and object masses, colors, sizes, etc. Different causal studies like do-interventions, counterfactual situations, structure learning, inference, etc. can be performed and evaluated using this platform.
Advection Diffusion:
The Advection Diffusion simulation (Ebert-Uphoff and Deng, 2015) mimics two important dynamic processes of many geoscience systems addiction and diffusion of a substance or property by a fluid or air. This is a spatiotemporal dataset of different properties of these processes in two grid sizes of 10x10 and 20x20. This dataset contains different scenarios ranging from simple to complex controlled by advection velocity field, advection and diffusion parameters, numerical parameters, spatial and temporal resolution, etc. parameters of these processes, and multiple runs of each scenario.
5.3 Real-world datasets
These are real observational datasets from the physical systems. So we don’t know the ground truth effect for these datasets, but we can generate counterfactuals from the knowledge learned from these datasets and measure future effects.
Beijing Multi-Site Air-Quality Dataset:
The Beijing Multi-Site Air-Quality Dataset (Chen, 2019) is an observational dataset collected by the Beijing Municipal Environmental Monitoring Center from different monitoring stations. This dataset contains hourly observation of 6 pollutants in the air: PM2.5, PM10, SO2, NO2, CO, and O3, and 6 meteorological variables: air temperature, wind direction and speed, pressure, dew point temperature, and precipitation. These data were collected from March 1st, 2013 to February 28th, 2017.
ERA5:
ERA5 (Hersbach et al., 2023) is a global climate and weather dataset maintained by the European Centre for Medium-Range Weather Forecasts (ECMWF). This dataset contains hourly observations of the different atmospheric, ocean-wave, and land-surface variables from 1940 to the present day. This dataset covers the observations from the whole globe and the repository is updated daily for new data.
Sea Ice Data:
The NSIDC DAAC SMMR and SSM/I-SSMIS data collection (DiGirolamo et al., 2022) is a polar sea ice observational dataset maintained by the National Snow and Ice Data Center (NSIDC). This dataset is collected from the Scanning Multichannel Microwave Radiometer (SMMR) instrument on the Nimbus-7 satellite and the Special Sensor Microwave/Imager (SSM/I) and Special Sensor. The data collection consists of observational measurements of sea ice concentration and extent, sea surface temperatures, wind stress, snow cover, rainfall rates, and more variables from 1978 to the present.
Metropolit Cohort:
The Metropolit Cohort dataset contains data from 11532 men born in 1953 in the Copenhagen Metropolitan area who were living in Denmark in 1968 (Osler et al., 2006). This is a longitudinal dataset of these men starting from their birth and data from different stages of their life is recorded. This dataset comprises physical, medical, mental, social, and diagnosis information collected from nationwide social and health registers. This is a high-quality dataset with good validity and a low degree of measurement error.
YahooFinancials:
Yahoo Financials  [^11] Yahoo Finance datasets. https://pypi.org/project/yahoofinancials/ is a data module consisting of financial data of daily stock prices of companies from the stock market. Using this dataset module we can get financial data for stock, cryptocurrency, forex, mutual fund, commodity futures, ETF, and US Treasury from Yahoo Finance for various time intervals and time ranges.
Lalonde:
The Lalonde dataset (LaLonde, 1986; Dehejia and Wahba, 1999) is a popular observational dataset collected from the National Supported Work Demonstration. The study looked at the effectiveness of a job training program (the treatment) on the real earnings of an individual, a couple of years after completion of the program. The dataset contains 445 observations, corresponding to 185 treated and 260 control subjects, a number of demographic variables (age, race, academic background, and previous real earnings), as well as a treatment indicator, and the real earnings in the year 1978 (the response). The treatment assignment indicator is the first variable of the data frame: treatment (1 = treated; 0 = control).
@@ 5.3 Real-world datasets
Besides the dataset discussed here, CauseMe (Muñoz-Marí et al., 2020) is a public platform that contains a large number of datasets with ground truth causal graphs to validate the causal discovery methods. This platform contains synthetic datasets with a wide range of properties like extreme events, complex dynamics, different errors, time delay, etc. resembling real dynamical systems. Real-world datasets with multiple dimensions are also included in this benchmarking platform where the causal graph is accepted by domain experts. It is also possible to submit new time series datasets from different domains to this platform.
@@ 5.4 Toolboxes for Causal Analysis
Table 9: Toolbox and APIs available for performing Causal Discovery and Causal Inference. Methods with (*) are suitable for time-series data.
Tools Methods Language Causal Discovery (CD), Causal Inference (CI)
pcalg (Markus Kalisch et al., 2012; Alain Hauser and Peter Bühlmann, 2012)
PC, stable PC, CPC, GES, GIES, ARGES, GDS, AGES, FCI, FCI–JCI, FCI+, RFCI, LINGAM, IDA algorithm, Generalized Backdoor Criterion (GBC), Generalized Adjustment Criterion (GAC)
R
CD, CI
Tetrad (Ramsey et al., 2018)
FCI, RFCI-BSC, FGES, GFCI, PC, PCStable, CPC, PcMax, RFCI, MBFS, GLASSO, FOFC, FTFC, LINGAM, TsFCI*, TsGFCI*, TsIMaGES*, MultiFASK*.
Java
CD
Causal-cmd (Bui et al., 2022)
BPC, EB, Fang-Concatenated, FAS, Fask-Concatenated, FCI, FGES, FGES-MB, FOFC, FTFC, GFCI, GLASSO, IMGS_CONT, IMGS_DISC, MBFS, MGM, PC-ALL, PC-STABLE-MAX, R-SKEW, R-SKEW-E, R1, R2, R3, R4, RFCI, SKEW, SKEW-E, TS-FCI*, TS-GFCI*, TS-IMGS*
Java
CD
causal-learn (Zheng et al., 2023)
PC, FCI, CD-NOD, GES, ICA-based LiNGAM, DirectLiNGAM, VAR-LiNGAM*, RCD, CAM-UV, Additive noise models, Generalized Independence Noise (GIN) condition-based method, GRaSP, Granger causality
Python
CD
CausalNex (Beaumont et al., 2021)
Bayesian Networks, Counterfactual analysis, Identify the right intervention
Python
CD, CI
LiNGAM (Ikeuchi et al., 2023)
DirectLiNGAM, MultiGroupDirectLiNGAM, VARLiNGAM*, VARMALiNGAM*, Bootstrap, RCD, RESIT, LiM, CAM-UV, MultiGroupRCD
Python
CD
Tigramite  [^12] https://github.com/jakobrunge/tigramite
PCMCI*, PCMCI+*, LPCMCI*, RPCMCI*, J-PCMCI+*
Python
CD
Causal ML (Chen et al., 2020)
Uplift Random Forests, Interaction Tree, Causal Inference Tree, S-Learner, T-Learner, X-Learner, R-Learner, Doubly Robust (DR) learner, 2-Stage Least Squares (2SLS), Doubly Robust Instrumental Variable (DRIV) learner, CEVAE, DragonNet
Python
CI
EconML (Keith et al., 2019)
Double Machine Learning (RLearner), Dynamic Double Machine Learning, Causal Forests, Orthogonal Random Forests, T-Learner, S-Learner, X-Learner, Domain Adaptation Learner, Doubly Robust Learner, Non-Parametric Double Machine Learning, Deep Instrumental Variables, Dynamic Double Machine Learning
Python
CI
DoWhy (Sharma and Kiciman, 2020; Blöbaum et al., 2022)
Average causal effect for backdoor, frontdoor, instrumental variable, and other identification methods
Python
CI
linearmodels (Sheppard et al., 2023)
Fixed effects, First difference regression, Between estimator for panel data, Pooled regression for panel data, Fama-MacBeth estimation of panel models, Absorbing Least Squares, Two-stage Least Squares, Limited Information Maximum Likelihood, k-class Estimators, Generalized Method of Moments, 2- and 3-step estimation, Time-series estimation, GMM estimation, Seemingly Unrelated Regression (SUR/SURE), Three-Stage Least Squares (3SLS), Generalized Method of Moments (GMM) System Estimation
Python
CI
Causal Discovery Toolbox (Kalainathan and Goudet, 2019)
CGNN, PC, GES, GIES, LiNGAM, CAM, GS, IAMB, MMPC, SAM, CCDr, ANM, IGCI, RCC, NCC, GNN, Bivariate fit, Jarfo, CDS, RECI
Python
CD, CI

### v2

@@ 5 Datasets and Toolbox for Causality
(Smith et al., 2011) Each of these causally sufficient dataset contains the neural activity based on the blood flow change and up to 50 time series Discovery Simulated
series with 50 and 5000 time steps.
DREAM4 Is a simulated gene expression dataset based on patterns found in model organisms. This dataset contains time series and Time Causal Realistic
(Marbach et al., 2009) steady-state settings data for learning the gene regulatory network. Five distinct datasets are available in DREAM4, series Discovery Simulated
each with ten distinct time-series recordings for one hundred genes spread over 21 time steps.
NCEP-DOE Reanalysis 2 The US National Centers for Environmental Protection (NCEP) and the Department of Energy (DOE) provide this Spatio- Causal Realistic
product dataset from 1979 to the current time. All available data is applied to a complex climate model to generate temporal Discovery Simulated
(Kanamitsu et al., 2002) reanalysis data for unobserved locations and missing time steps. This is a large set of almost 40 atmospheric variables
measured in the reanalysis dataset. The dataset covers 90N-90S, 0E-357.5E with a
2.5-degree latitude x 2.5-degree longitude global grid (144x73).
Table 10: (Continued) Synthetic, simulated, and real world datasets used for causal analysis.
Dataset Description Data Causal Dataset
Type Category Type
CausalWorld This open-source causal structure learning benchmarking data generation platform contains a robotic environment Time-series Causal Realistic
manipulation dataset for different tasks. The generated datasets represent different causal structures of interacting Discovery / Simulated
objects like robot and object masses, colors, sizes, etc. Different causal studies like do-interventions, counterfactual Inference
situations, structure learning, inference, etc. can be performed and evaluated using this platform.
Beijing Multi-Site This observational dataset was collected by the Beijing Municipal Environmental Monitoring Center and contains Time-series Causal Real-world
Air-Quality Dataset hourly observation of 6 pollutants in the air: CO, PM2.5, PM10, O3, NO2 and SO2, and 6 meteorological variables: Discovery /
(Chen, 2019) air temperature, wind direction and speed, pressure, dew point temperature, and precipitation. These data were Inference
collected from 2013 to 2017.
ERA5 The European Centre for Medium-Range Weather Forecasts (ECMWF) maintains this global climate and weather Spatio- Causal Real-world
(Hersbach et al., 2023) dataset. The hourly observations of the different atmosphere, land and oceanic variables are available temporal Discovery /
in this dataset from 1940 to the present day for the whole globe and are updated daily for new data. Inference
Sea Ice Data This data collection is a polar sea ice observational dataset maintained by the National Snow and Ice Data Spatio- Causal Real-world
(DiGirolamo et al., 2022) Center (NSIDC). This dataset is collected from the Scanning Multichannel Microwave Radiometer temporal Discovery /
(SMMR) instrument on the Nimbus-7 satellite and the Special Sensor Microwave/Imager (SSM/I). Several Inference
observational variables like sea ice concentration and extent, sea surface temperatures, wind stress, snow
cover, rainfall rates, etc. are recorded in this dataset from 1978 to the present.
Metropolit Cohort The Metropolit Cohort dataset contains data from 11532 humans born in 1953 and lived till 1968 in the Copenhagen Time-series Causal Real-world
(Osler et al., 2006) Metropolitan area, Denmark. This dataset comprises physical, medical, mental, social and diagnosis Discovery
information from different stages of life of these men collected from nationwide social and health registers.
This is a very reliable dataset with minimal measurement error and strong validity.
YahooFinancials [^12] https://pypi.org/project/yahoofinancials/ This data module consists of financial data of daily stock prices of companies from the stock market. Tim-series Causal Real-world
Using this dataset module we can get financial data of different transactions from Yahoo Finance like stock, Inference
mutual fund, cryptocurrency, ETF, US Treasury and forex for various time intervals and time ranges.
Lalonde This is a popular observational dataset collected from the National Supported Work Demonstration. The study Time-series Causal Real-world
(LaLonde, 1986) examined how well a work training program (the treatment) affected a participant’s actual wages a few years Inference
Dehejia and Wahba (1999) after the program’s conclusion. Besides the treatment indicator the dataset provides demographic variables
like age, race, academic background and previous real earnings for 260 controlled and 185 treated subjects
(a total of 445) with the response (real earnings in the year 1978). In the values of the treatment assignment
indicator variable 1 means treated and 0 means control/untreated.
@@
Besides the datasets discussed here, CauseMe (Muñoz-Marí et al., 2020) is a public platform containing numerous ground truth causal graphs and corresponding datasets to validate the causal discovery methods. This platform contains synthetic datasets with a wide range of properties like extreme events, complex dynamics, different errors, time delay, etc. resembling real dynamical systems. Real-world datasets with multiple dimensions are also included in this benchmarking platform where the causal graph is accepted by domain experts. It is also possible to submit new time series datasets from different domains to this platform.
@@ 5.1 Toolboxes for Causal Analysis
Table 11: Tools for performing Causal Discovery (CD) and Causal Inference (CI). Methods with (*) are suitable for time-series data.
Tools Methods Language Category: CD / CI
pcalg (Markus Kalisch et al., 2012) PC, stable PC, CPC, GES, GIES, ARGES, GDS, AGES, LINGAM, FCI, FCI–JCI, R CD, CI
(Alain Hauser and Peter Bühlmann, 2012) FCI+, RFCI, Generalized Adjustment Criterion (GAC),
IDA algorithm, Generalized Backdoor Criterion (GBC)
Tetrad (Ramsey et al., 2018) FCI, RFCI-BSC, FGES, GFCI, PC, PCStable, CPC, PcMax, RFCI, MBFS, GLASSO, FOFC, FTFC, Java CD
LINGAM, TsFCI*, TsGFCI*, TsIMaGES*, MultiFASK*.
Causal-cmd (Bui et al., 2022) FCI, FGES, Fask-Concatenated, FGES-MB, BPC, EB, Fang-Concatenated, FTFC, FAS, FOFC, Java CD
GFCI, MGM, PC-ALL, PC-STABLE-MAX, GLASSO, MBFS, IMGS_CONT, IMGS_DISC,
RFCI, R-SKEW, R-SKEW-E, SKEW, SKEW-E, R1, R2, R3, R4, TS-FCI*, TS-GFCI*, TS-IMGS*
causal-learn (Zheng et al., 2023) PC, FCI, CD-NOD, GES, ICA-based LiNGAM, DirectLiNGAM, VAR-LiNGAM* Python CD
RCD, CAM-UV, Additive noise models, Generalized Independence Noise (GIN)
condition-based method, GRaSP, Granger causality
CausalNex (Beaumont et al., 2021) Bayesian Networks, Counterfactual analysis, Identify the right intervention Python CD, CI
LiNGAM (Ikeuchi et al., 2023) DirectLiNGAM, MultiGroupDirectLiNGAM, VARLiNGAM*, VARMALiNGAM*, Python CD
Bootstrap, RCD, RESIT, LiM, CAM-UV, MultiGroupRCD
Tigramite  [^13] https://github.com/jakobrunge/tigramite PCMCI*, PCMCI+*, LPCMCI*, RPCMCI*, J-PCMCI+* Python CD
Causal ML (Chen et al., 2020) Interaction Tree, Causal Inference Tree, Uplift Random Forests, S-Learner, T-Learner, Python CI
R-Learner, X-Learner, Doubly Robust (DR) learner, 2-Stage Least Squares (2SLS),
Doubly Robust Instrumental Variable (DRIV) learner, DragonNet, CEVAE
EconML (Keith et al., 2019) Causal Forests, Orthogonal Random Forests, S-Learner, T-Learner, X-Learner, Python CI
Double Machine Learning (RLearner), Dynamic Double Machine Learning, Domain Adaptation Learner,
Doubly Robust Learner, Non-Parametric Double Machine Learning, Deep Instrumental Variables,
Dynamic Double Machine Learning
DoWhy (Sharma and Kiciman, 2020) Average causal effect for backdoor, frontdoor, Python CI
(Blöbaum et al., 2022) instrumental variable, and other identification methods
linearmodels (Sheppard et al., 2023) Fixed effects, Between estimator for panel data, First difference regression, Pooled regression for Python CI
panel data, Fama-MacBeth estimation of panel models, Absorbing Least Squares, Two-stage Least
Squares, k-class Estimators, Limited Information Maximum Likelihood, Generalized Method of
Moments, GMM estimation, Time-series estimation, 2- and 3-step estimation, Seemingly Unrelated
Regression (SUR/SURE), Three-Stage Least Squares (3SLS), Generalized Method of Moments (GMM)
Causal Discovery Toolbox CGNN, PC, GES, GIES, LiNGAM, CAM, GS, IAMB, MMPC, SAM, CCDr Python CD, CI
(Kalainathan and Goudet, 2019) ANM, IGCI, RCC, NCC, GNN, Bivariate fit, Jarfo, CDS, RECI